OpenStudy (anonymous):

which ordered pair is a solution of the equation y = -7x + 2 (1,2) (8,-54) (5,-35) (1,-7) i think its d. but im not sure.

OpenStudy (nfcfox):

It's not d. Try testing all the coordinates. Again just take the x's and plug it in and solve for y, if the y is the same as the y in the coordinate then that's your answer.

OpenStudy (anonymous):

c? im bad at these.

OpenStudy (nfcfox):

No, b. If we plug in 8 we will get y= -7(8)+2. Which is also y= -56+2 y= -54. Go back to our coordinate and see the y value is -54 when x is 8. So our answer is B.